# Identifying noise transients in gravitational-wave data arising from nonlinear couplings

### v1

@@ 5 Conclusion
It is important to stress that introducing artificial phase randomization in analyzing the data does not affect the actual phase coupling in the signal, but reduces the false detection of such couplings. We demonstrated this with simulations as well as an example of a glitch that appeared in real GW data. We used the PR auto-bicoherence to examine the similarities and differences in the phase-coupled frequencies within certain categories of glitches as well as across them. This third-order statistic provides an additional way of characterizing glitches than Omicron spectrograms and may contribute to troubleshooting their origins and subsequent removal or mitigation. It is also important to realize that the bicoherence tool developed here is not limited to characterizing glitches and can identify other features in GW data, such as the interesting frequency evolution of CBC signals, even though this is not its primary objective. We demonstrated this with a couple of examples of actual GW events and simulated signals as proof of concept.
@@ Appendix B Similar Bicoherence Morphology
A few examples of glitches with similar morphology in the third-order statistic are shown in Fig. 9. The top row there displays the time-frequency maps of those glitches and is classified by Gravity Spy into the “extremely loud” and “koi fish” categories. The bottom row shows the corresponding bicoherence maps.
@@
Appendix C Cross-bicoherence: A feasibility study

### v2

@@ 5 Conclusion
It is important to stress that introducing artificial phase randomization in analyzing the data does not affect the actual phase coupling in the signal, but reduces the false detection of such couplings. We demonstrated this with simulations as well as an example of a glitch that appeared in real GW data. We used the PR auto-bicoherence to examine the similarities and differences in the phase-coupled frequencies within certain categories of glitches as well as across them. This third-order statistic provides an additional way of characterizing glitches than Omicron spectrograms and may contribute to troubleshooting their origins and subsequent removal or mitigation.
@@
In addition to checking if certain noise transients, e.g., flagged by Omicron [38] or GravitySpy [36], have origins in nonlinear coupling, bicoherence calculations can complement machine learning studies like DeepClean and NonSENS by confirming the presence of phase coupling as well as provide more information about the nature of the nonlinear coupling (such as the specific frequencies and channels that are coupled), which can be useful in regressing that excess noise in software and, in some cases, through hardware fixes. Conversely, phase couplings found first by bicoherence may be used for training machine learning algorithms to aid them in regressing excess GW strain noise arising from them. Whether other noise modeling and regression tools such as BayesWave [16] and SHAPES [21] can benefit from bicoherence computations remains to be seen. We will report on these investigations in a future work [34].
@@ Appendix B Similar bicoherence morphology
A few examples of glitches with similar morphology in the third-order statistic are shown in Fig. 8. The top row there displays the time-frequency maps of those glitches and is classified by Gravity Spy [36] into the “extremely loud” and “koi fish” categories. The bottom row shows the corresponding bicoherence maps.
@@
Appendix C Bicoherence maps of some short duration compact binary coalescence signals
By applying the PR auto--bicoherence analysis to GWTC-3 [7] catalog events we examine the apparent phase coherence in the data with CBC signals. We noticed significant bicoherence points in a few GW events. Two of those, namely, GW events GW150914 [57], and GW190521_074359 [5] are presented in Fig. 9 along with a simulated signal for aiding the interpretation of their bicoherence maps. The presence of significant bicoherence is evident there. To prepare the GW strain data for the bicoherence analysis, first a high-pass filter is applied with cut-off frequency 20 Hz, followed by whitening and the application of a bandpass filter, from 30 to 350 Hz, with notches of the first three power line harmonics [58]. The noise power-spectral density (PSD) is estimated from 32 sec of data using the Welch method.
To ensure that the phase couplings revealed in GW events’ bicoherence maps are not from glitches, we modeled two signals similar to the GW150914 event, with and without strain noise. In Fig. 9 (row 1, column 3), the 8 second long time-series noise was simulated by using the H1 noise PSD available in Ref. [59] for the GW190514 event. A simulated CBC signal was added to that noise before producing the plot. On the other hand, in Fig. 9 (row 2, column 3) no noise was added to the CBC signal. The bicoherence patterns in the map of the modeled signal, with added simulated noise, resembles those of the GW150914 signal, as shown in Fig. 9 (row 1, column 3). These signals were produced from the SEOBNRv4_opt waveform [60], with component masses of 36 Mdirect-$\odot$ and 30 Mdirect-$\odot$. This family of waveforms was constructed in the literature by calibrating effective one-body waveforms with an extensive set of numerical relativity waveforms, and have been employed in LIGO-Virgo CBC searches (see, e.g., Ref. [61]). The matched-filter SNR of the signal is 20. At the very least, the knowledge of these bicoherence patterns will be important in not confusing real CBC signals for noise artifacts and, therefore, for the purpose of veto safety. We exphasize that bicoherence maps are not the optimal method for detecting such CBC signals; rather the matched-filtering technique is [62].
Appendix D Cross-bicoherence: A feasibility study